A random sample of eight drugstores shows the following prices (in $ of a popular pain reliever:
3.50, 4.00, 2.00, 3.00, 2.50, 3.50, 2.50, 3.00
2.50 3.00
a. What is the standard deviation for the sample?
b.Assume the normal distribution for the underlying population to construct the 90% confidence interval for the population mean.
1
Expert's answer
2020-07-13T19:00:25-0400

s2 = Σ(xi - x̄)2/(N - 1)=3.225/9= 0.35833

s = 0.5986

For confidence 90%: 1.645s

Using table margin 2.95 ±0.311 (±10.56%)